OpenStudy (anonymous):

write the equation of the line that is perpendicular to the line y=-3x+4 that goes through the point (6,10).

OpenStudy (jdoe0001):

hmmm what's the slope of say " y=-3x+4 " ?

OpenStudy (jdoe0001):

? so... the slope of that one is..... *rolling drums* ?

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

you may want to recheck your book on the slope-intercept form and brush up a bit on it I'd think

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

I got y=-1/3x+12

OpenStudy (michele_laino):

two lines are perpendicular each other, when the subsequent relationship, holds: \[\large m \times m' = - 1\] where m and m' are the slopes of those lines

OpenStudy (michele_laino):

what is the slope of the perpendicular line?

OpenStudy (anonymous):

y=-1/3x+12

OpenStudy (anonymous):

final answer

OpenStudy (anonymous):

"write the equation of the line that is perpendicular to the line y=-3x+4 that goes through the point (6,10)."

OpenStudy (michele_laino):

I can not give the answer directly, since the Code of Conduct

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

hint: \[\Large \left( { - 3} \right) \times m' = - 1\] what is m'=...?

OpenStudy (anonymous):

it is 1/3

OpenStudy (anonymous):

sorry

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

no worries :)

OpenStudy (michele_laino):

now, the requested equation, is: \[\Large y - 10 = \frac{1}{3}\left( {x - 6} \right)\] please simplify that equation

OpenStudy (anonymous):

y=13x-68

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{3}\left( {x - 6} \right) = \frac{x}{3} - 2\]

OpenStudy (michele_laino):

so we have: \[\Large y - 10 = \frac{x}{3} - 2\]

OpenStudy (michele_laino):

now I add 10 to bot sides, so I get: